OpenStudy (erica399):

Can anyone help me with trigonometric identities? (sum and difference identities, evaluating expressions, common angles, Pythagorean identity)

OpenStudy (subha):

Tell your question.

OpenStudy (erica399):

First, I am supposed to apply the definitions of the sum and difference identities for cosine to 1/2(cos(a-b)-cos(a+b))

OpenStudy (erica399):

And then simplify

OpenStudy (subha):

Cos(a-b)=cos a cos b -sin a sin b

OpenStudy (erica399):

Right, and then it needs to be cos(a+b)=cosacosb-sinasinb, right?

OpenStudy (subha):

Similarly, cos(a+b)=cos a cos b - sin a sin b

OpenStudy (subha):

sorry,

OpenStudy (erica399):

Okay, from there, how do I simplify or combine the two? And what happened to the (1/2)?

OpenStudy (subha):

cos(a-b)=cosa cosb +sina sinb

OpenStudy (erica399):

I am confused, could you explain?

OpenStudy (subha):

now, this simplifies to 1/(2sin a sin b)

OpenStudy (subha):

you understand

OpenStudy (subha):

1/4sina sinb

OpenStudy (erica399):

Yea! Now, the other side of the equation (which I haven't given you yet) shows sin(a)sin(b), so now I have sin(a)sin(b)=1/(2sin(a)sin(b))

OpenStudy (erica399):

right?

OpenStudy (subha):

sin(a)sin(b)=1/(2sin(a)sin(b))

OpenStudy (subha):

this expression is right?

OpenStudy (erica399):

Well originally the equation was sinasinb=1/2(cos(a-b)-cos(a+b) Then we used the definitions of the sum and difference identities to get cos(a-b)=cosacosb+sinasinb and cos(a+b)=cosacosb-sinasinb Then we simplified to 1/(2sinasinb) So now it should be sinasinb=1/(2sinasinb)

OpenStudy (subha):

i think it is given in this way: sina sinb =(1/2)(cos(a-b)-cos(a+b))

OpenStudy (erica399):

Originally, yes

OpenStudy (subha):

actually cos terms are in numerator

OpenStudy (subha):

Am i right

OpenStudy (subha):

cos(a-b)-cos(a+b)=2sin a sin b (1/2)*2 sin a sin b =sin a sin b

OpenStudy (subha):

I think you get your answer.

OpenStudy (erica399):

Thank you @subha